A long straight wire carries a current 1.5A. an electron travels with a speed of 5*10^6cm/sec parallel to the wire 10cm from it and in the same direction as the current. What force does the magnetic field of the current exert on the moving electron
1
Expert's answer
2020-05-19T10:39:54-0400

First of all let's see the magnitude of magnetic field created by wire carrying current

"B=\frac{\mu_oI}{2\pi r}=\frac{4\pi\times10^{-4}\times1.5}{2\pi\times0.1}=3\times10^{-3}T"


force due to magnetic field ="q(v\times B)=1.6\times10^{-19}(5\times10^6\times3\times10^{-3})=2.4\times10^{-15}N" =